THE PURE INVERSION-TUNNELING TRANSITION OF AMMONIA IN HELIUM DROPLETS

\underline{WOLFGANG JÄGER} AND RUDOLF LEHNIG, Department of Chemistry, University of Alberta, Edmonton, AB, T6G 2G2.

The pure inversion tunneling transition of ammonia embedded in superfluid helium droplets has been measured for the two isotopomers ^{14}NH_{3} and ^{15}NH_{3}. The transitions were found at 21.72 GHz for ^{14}NH_{3} and at 20.68 GHz for ^{15}NH_{3} and are therefore shifted by 1.97 GHz and 1.94 GHz to the red compared to the gas phase values of the two isotopomers, respectively [Pickett et al., J. Quant. Spectrosc. Radiat. Transfer 60, 883 (1998).]. The measured peaks consist of a broad background with a width at half maximum of 1.5 GHz and an unexpectedly sharp peak with a width at half maximum of only 15 MHz sitting on top of the broad background. The difference between gas phase and droplet spectra will be discussed and a interpretation of the complex line shape will be given.